You wish to grow a thin film of anthracene using thermal evaporation. If your thermal evaporation chamber can reach pressures as low as 10.0 mPa, to what temperature do you need to heat anthracene to cause it to sublime?


1
Expert's answer
2021-05-18T06:56:03-0400

Pressure = P1 = 10 × 10-3 pa .

T1 = 298 k .


Temperature of Sublimation = T2 = ?

Pressure = P2 = 105 pa




P1 / T1 = P2 /T2



T2 = P2 × T1 / P1 .


T2 = 105 × 298 / 10-3 .


T2 = 298 × 108 K .